can someone please help me?
the perimeter of a rectangle field is 110 feet. the length is 7 feet more than twice the width. what is length of the rectangle?

Answer by nic0le116(25)

What can we use to represent the value of the width at this time?
Since we have no numerical value at this time, we'll use "x."
And what can we use for the length at this time?
Well we know that the length is 7ft longer than 2 times the width.
So we can represent this by 2x+7.
----------------------------------
We know that perimeter is the sum of all the sides.
In the case of the rectangle we can assume that perimeter equals 2 times the length+ 2 times the width.
----------------------------------
So with this information, let's set up an equation to solve for x
2x + 2(2x+7)=110 <--------- 2WIDTH+2LENGTH=PERIMETER
--------------------------------------
Solve for x!
2x+2(2x+7)=110 (Distribute the 2)
2x+4x+14=110 (Combine like terms)
6x+14=110 (Subtract 14 from each side)
6x=96 (Divide each side by 6)
x=16 <-------What does this x represent? The value of the width!
---------------------------------------
Now remember to always check what specifically the word problem is asking of you.
We want to know the value of the length, so we're not done yet.
So now we need to plug in the value of x into what we originally represented length to equal which was (2x+7)
------------------------------------
2(16)+7
32+7
39<------------------the value of the length!
---------------------------------------
Not sure if this is right? Check by adding up all the side of the values of width and length and see if they equal the given perimeter value.
39+39+16+16=110 CHECK!

ANSWER=
LENGTH=39ft